OpenStudy (xenesis98):

Alex stocks up for winter He buys 36 cans of vegetables. he pays 80 cents per can for tomatoes and 40 cents per can for corn for a total cost of $18. how many cans of corn does he buy?

OpenStudy (xenesis98):

just need the answer

OpenStudy (texaschic101):

t + c = 36 (tomatoes + corn = 36 cans .80t + .40c = 18 (80 cents per can for tomatoes + 40 cents Per can for corn = 18) those are the 2 equations needed to solve this problem. We can now solve by substitution. t + c = 36 t = 36 - c now sub 36 - c in for t in the other equation. .80t + .40c = 18 .80(36 - c) + .40c = 18 now, can you finish and solve for c ??